OpenStudy (anonymous):

I will fan and medal can you please help mw with the steps Part A: Factor 3x^2y^2 − 2xy^2 − 8y^2. Show your work. (4 points) Part B: Factor x^2 + 10x + 25. Show your work. (3 points) Part C: Factor x^2 − 36. Show your work. (3 points) these are my answers part a: y^2=(3x+4)(x-2) part b: (x+5)(x+5) part c: (x+6)(x-6)

OpenStudy (jdoe0001):

so....have you done any factoring? have you covered factoring?

OpenStudy (anonymous):

yes I have

OpenStudy (anonymous):

I can check my answer using the FOIL method I just don't know the steps to factoring if that makes sense @jdoe0001

OpenStudy (jdoe0001):

ok well.. let's start backwards any ideas on say C) since is very simple \(\bf x^2-36\) ?

OpenStudy (anonymous):

are the answers I put at the bottom of my question right @jdoe0001

OpenStudy (jdoe0001):

hmm ok well, those are given, but anyhow \(\bf x^2-36\qquad {\color{brown}{ 6^2=36}}\qquad thus\implies x^2-6^2 \\ \quad \\ --------------------------\\ \textit{difference of squares} \\ \quad \\ (a-b)(a+b) = a^2-b^2\qquad \qquad a^2-b^2 = (a-b)(a+b)\\ --------------------------------------\\ x^2-6^2\implies (x-6)(x+6)\)

OpenStudy (anonymous):

okay so I would explain this one by saying the square root of 36 is 6

OpenStudy (jdoe0001):

yes, or that 36 is the square of 6 and thus it makes it a difference of squares

OpenStudy (anonymous):

so that is showing work on factoring?

OpenStudy (jdoe0001):

for C), yes

OpenStudy (jdoe0001):

B) is also simple if we factor 25, we get 5, 5, 1 5*5*1 = 25 if we "multiply" then 5*5 we get the last term if we "add" 5+5, we get the middle term thus \(\begin{array}{cccllll} x^2&+10&+25\\ &\uparrow &\uparrow \\ &5+5&5\cdot 5 \end{array}\implies (x+5)(x+5)\)

OpenStudy (anonymous):

okay so if I were to explain it, it would be in order to factor x^2+10x+25 you pull out the GCF which in this case is 5 and simplify to (x+5)(x+5)

OpenStudy (jdoe0001):

well... 5 is not the GCF, notice \(x^2\) doesn't have a 5 you'd factor out the last term, the constant one and from a combination of two of its factors you ought to get the middle term

OpenStudy (anonymous):

so from the original problem of x^2+10x+25 you would factor the 10 and 25 and that's is supposed to be explaining the steps of how you get this answer?

OpenStudy (jdoe0001):

hmm well... you said you have already covered factoring so...... you should recall that you need to factor out the constant to get the values for the binomials

OpenStudy (triciaal):